OpenStudy (anonymous):

NO IDEA HOW TO SOLVE THIS.... Solve the following system of equations. 2a - 3b + c = 102a - 2b - 2c = 2a + 3b + 2c = -1 (2, 3, 4) (-1, 2, 3) (1, -2, 2) (3, -1, 0)

OpenStudy (mathstudent55):

Three equations in three variables: 2a - 3b + c = 102 a - 2b - 2c = 2 a + 3b + 2c = -1 is this the correct system of equations?

OpenStudy (anonymous):

yes sir

OpenStudy (anonymous):

same question,

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

ok, if you add all the equations, you get

OpenStudy (anonymous):

4a - 2b + c = 103

OpenStudy (anonymous):

use substitution method to find what c is

OpenStudy (mathstudent55):

Ok, you need to take two equations and add them together or subtract and eliminate one variable. Then take another 2 equations and eliminate the same variable. Let's eliminate c. First add eq 2 and eq 3: 2a + b = 1 here's a new eq with no c

OpenStudy (mathstudent55):

Now if you add eq 1 and eq 2, c will not eliminate, so first multiply the entire eq 1 by 2: 4a - 6b + 2c = 204 and add to eq 2 a - 2b - 2c = 2 --------------------- 5a - 8b = 206

OpenStudy (anonymous):

The answer isn't C ;)

OpenStudy (mathstudent55):

This is another eq without c. Now solve the two equations without c as a system of 2 equations in two variables

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

-1, 2, 3

OpenStudy (mathstudent55):

I think the system of eqs is incorrect. Your question doesn't have spaces between the equations. I think I wrote the 2nd and 3rd equations wrong.

OpenStudy (anonymous):

Don't worry, you're still smart as h***

OpenStudy (mathstudent55):

I figured out what the correct problem was. I wrote it out and scanned it. See pdf above.

OpenStudy (anonymous):

youre the BEST

OpenStudy (mathstudent55):

thanks

OpenStudy (anonymous):

no thank you, really.